What equation should I use? the area of a rectangular room is 26 square meters.
The long dimension of the room is 5 meters longer than the short dimension.

Answer by mananth(16946)

let width be x m
length = x+5 m
..
Area = L*W
x*(x+5)=26
x^2+5x=26
x^2+5x-26=0
use quadratic equation
discriminant = 129
x1= (-5+sqrt(129))/2
x1=3.18
x2 =(-5-sqrt(129))/2
x2 will be negative so ignore
width = 3.18
length = 3.18 +5 = 8.18
...
check
8.18*3.18=26 m^2
...
m.ananth@hotmail.ca